OpenStudy (froggirl):

Tina’s team entered a math contest where teams of students compete by answering questions that are worth either 3 or 5 points. No partial credit is given. Tina’s team scored 44 points on questions. How many 5-point questions did the team answer correctly?

OpenStudy (amistre64):

3x+5y = 44 this might be applicable to number theory

OpenStudy (amistre64):

since the gcd of 3 and 5 is 1 determine the solutions to: 3x+5y = 1 5 = 3(1) + 2 3 = 2(1) + 1 0 0 1 0-1(1) = -1 1-1(-1) = 2 ; x=2 3(2)+5(-1) = 1 ; y = -1 times it by 44 3(88)+5(-44) = 44 x = 88 + 5n y = -44 - 3n solve n for x,y >0 88 + 5n > 0 ; n > -88/5, -17,-16,-15 .... -44 - 3n >0 ; n < -44/3 -14,-15,-16 looks like -14 and -15 might be useful if i remembered it correctly

OpenStudy (amistre64):

i think i erred that 44 = 5(8) + 4 44 = 5(7) + 9 <-- possible 3(3) 44 = 5(6) + 14 44 = 5(5) + 19 44 = 5(4) + 24 <-- possible 3(8) 44 = 5(3) + 29 44 = 5(2) + 34 44 = 5(1) + 39 <-- possible 3(13) 44 = 5(0) + 44